OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 64 cm3? @Pawtpie

OpenStudy (anonymous):

I'll try to keep my responses shorter lol

OpenStudy (anonymous):

V=s^3

OpenStudy (anonymous):

no it is fine this is the whole ?

OpenStudy (anonymous):

64=s^3 solve for s

OpenStudy (anonymous):

Where V is volume and s is side length

OpenStudy (anonymous):

@shelbydancer ???

OpenStudy (anonymous):

sorry I needed a brake

OpenStudy (anonymous):

64/6=10.6666

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

The volume of a cube is the side length cubed

OpenStudy (anonymous):

So we need the cube root of the volume to get the side length

OpenStudy (anonymous):

more

OpenStudy (anonymous):

s^3 means you multiply s by s by s, so you have to take the cuberoot of 64, so you have to ask yourself what do you multiply by itself 3 times to get 64? For example 2 times 2 times 2=8 so 2 is the cbrt of 8

OpenStudy (anonymous):

multiply s by s by s is that a 5 or a s

OpenStudy (anonymous):

s, the side length of the cube

OpenStudy (anonymous):

s, and s stands for side length, what you're trying to solve for

OpenStudy (anonymous):

so how do I do this???

OpenStudy (anonymous):

So we need to find what number, when multiplied by itself 3 times equals 64.

OpenStudy (anonymous):

64/3=

OpenStudy (anonymous):

21

OpenStudy (anonymous):

\[s * s * s = 64cm^3\]

OpenStudy (anonymous):

And we need to find s

OpenStudy (anonymous):

Do you know how square rooting works first of all?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Ok for instance, lets find the square root of 9.

OpenStudy (anonymous):

That means, what number, when multiplied by itself, gives us 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct, 3 times itself, so 3 * 3 = 9

OpenStudy (anonymous):

So 3 is the square root of 9

OpenStudy (anonymous):

What we need to do is find the CUBE root of 64

OpenStudy (anonymous):

So a number, multiplied by itself, multiplied by itself again, equaling 64

OpenStudy (anonymous):

Lets just take a random number

OpenStudy (anonymous):

What would be 3 cubed \[3^3\]

OpenStudy (anonymous):

So \[3 * 3 * 3\]

OpenStudy (anonymous):

27

OpenStudy (anonymous):

Okay, so that means 3 is the cube root of 27

OpenStudy (anonymous):

side of a cube with a volume of 64 cm3?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So a number times itself 3 times equaling 64

OpenStudy (anonymous):

please hurry my mom wants me done tonight by 11:30

OpenStudy (anonymous):

Try 4

OpenStudy (anonymous):

so 4 * 4 * 4 is

OpenStudy (anonymous):

64

OpenStudy (anonymous):

So the side length is 4 cm.

OpenStudy (anonymous):

thax

OpenStudy (anonymous):

What time is it currently?

OpenStudy (anonymous):

It's 1:06 AM here

OpenStudy (anonymous):

where are u

OpenStudy (anonymous):

Eastern time

OpenStudy (anonymous):

o standered mt.

OpenStudy (anonymous):

How many questions are left?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

4 more 5 total

OpenStudy (anonymous):

Okay lets start them up!